OpenStudy (jennychan12):

Can someone show me how to use Simpson's Rule in integral (1/(x+1)^2)dx from 1 to 3?

OpenStudy (jennychan12):

\[\int\limits_{1}^{3} \frac{ x }{(x+1)^2 }dx\]

OpenStudy (kainui):

Never heard of it before.

OpenStudy (anonymous):

Ok, here's Simpson's Rule: \[\int\limits_a^b f(x) dx \approx \frac{b-a}{6}(f(a)+4f(\frac{a+b}{2})+f(b))\] So in your case: \[\int\limits_1^3\frac{x}{(x+1)^2} dx \approx \frac{1}{3}(0.25+\frac{8}{9}+\frac{3}{16} )\] Plug and chug.

OpenStudy (jennychan12):

My teacher never taught Simpson's rule, so I looked at my textbook and I found the formula but since the formula is 

OpenStudy (anonymous):

Oh, that's the composite Simpson's Rule. It goes 2, 4, 2, 4, etc.

OpenStudy (jennychan12):

it's n = 4

OpenStudy (jennychan12):

so it just repeats 2,4,2,4,2,4....until x(n) term right?

OpenStudy (anonymous):

Right, so for n = 4, it would be 1, 2, 4, 1

OpenStudy (jennychan12):

ok thanks. that's all i needed. :)

OpenStudy (anonymous):

Great, have fun integrating :P

OpenStudy (jennychan12):

haha thanks :D